OpenStudy (anonymous):

Find the exact values of the six trigometric functions of theta if theta is in standard positiion and the terminal side of theta is quadrant 4 and is on the line y=-3/7x

OpenStudy (mertsj):



OpenStudy (anonymous):

k then this would be square root 58 and then how would i ifind the 6 trigometric functions?

OpenStudy (mertsj):

\[\sin x=\frac{side opp}{hypotenuse}\]

OpenStudy (mertsj):

Use the definitions.

OpenStudy (anonymous):

but why do we use sin?

OpenStudy (mertsj):

What does the problem ask you to find?

OpenStudy (anonymous):

Find the exact values of the six trigometric functions of theta if theta is in standard positiion and the terminal side of theta is quadrant 4 and is on the line y=-3/7x

OpenStudy (mertsj):

So, the problem says to find the values of the 6 trig functions. What are the 6 trig functions?

OpenStudy (anonymous):

this is what i got sin theta =y/r=-3/sqrt 58 cos theta=x/r=7/sqrt58 tan=y/x=-3/7 csc theta=r/y=sqrt 58/3 sec=r/x=sqrt58/7 cot=x/y =-7/3

OpenStudy (mertsj):

csc should be negative. You should rationalize the denominators

OpenStudy (anonymous):

\[\sqrt{58\frac{ }{ ? }} \times\frac{ 3 }{ 3 }\]

OpenStudy (anonymous):

sqrt58 divided by 3

OpenStudy (mertsj):

sqrt58/3

OpenStudy (mertsj):

Denominator is already rational.

OpenStudy (anonymous):

i dont get why u say its a negative(csc theta).sorry for all the questions im trying to understand

OpenStudy (mertsj):

Because the csc is the reciprocal of the sin and the sin (which is like y) is negative in Quadrant IV

OpenStudy (anonymous):

so its sqrt 58 divided by -3?